OpenStudy (anonymous):

Solve. |y-1/4| = 6 (1/4 is a fraction)

OpenStudy (anonymous):

@mangorox ? any help?

OpenStudy (anonymous):

y-(1/4) = 6 // - 6 y-(1/4)-6 = 0 y-1/4-6 = 0 y-25/4 = 0 // + 25/4 y = 25/4 y = 25/4

OpenStudy (anonymous):

omg thank you so much! :) @mangorox